You and a friend are 4.5kilometers away from each other. Walking toward each other, the two of you meet in a half hour. Your walking speed is twice as fast as your friends. What is your friends walking speed?
Answer by lwsshak3(11628)

You and a friend are 4.5kilometers away from each other. Walking toward each other, the two of you meet in a half hour. Your walking speed is twice as fast as your friends. What is your friends walking speed?
let x=friend's walking speed
2x=your walking speed
distance=speed*time of travel
time of travel=1/2 hour
distance =4.5 kilometers
(1/2)x+(1/2)2x=4.5
(3/2)x=4.5
x=9/3=3
ans:your friend's walking speed is 3km/hr